The number of work-related injuries per month in a manufacturing plant is known to follow a Poisson distribution with a mean of 2.5 work-related injuries a month. What is the probability that in a given month no work-related injuries occur? What is the probability that atleast one work-related injury occurs?




1
Expert's answer
2021-01-26T04:06:45-0500

"P(x=0)=\frac{e^{-2.5}2.5^{0}}{0!}\approx 0.082\\

 P(x\geq 1)=1-P(x=0)\\
=1-\frac{e^{-2.5}2.5^{0}}{0!}=0.918"